OpenStudy (anonymous):

V(r)= - a e^2 lambda exp(-r/p) ---------------- + 4 pi epilson r

OpenStudy (anonymous):

show that force can be expressed as F(r) = a e^2 (-(r_0)^2/r) ------------------- 4 pi epilson r_0^2 + exp(- r -r_0/ rho))

OpenStudy (anonymous):

@Jemurray3

OpenStudy (anonymous):

so do I just take derivative

OpenStudy (anonymous):

erm

OpenStudy (anonymous):

\[ V(r) = \frac{-ae^2}{4\pi \epsilon_0 r^2} + \lambda e^{-r/p} \]?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

Yes, \[ \vec{F} = -\vec{\nabla} V \]

OpenStudy (anonymous):

In your case, since the potential is a function only of r, \[ \vec{F} = -\frac{dV}{dr} \hat{r} \]

OpenStudy (anonymous):

so , I get \[\frac{a e^2}{2 \pi r^3 \epsilon }-\frac{e^{-\frac{r}{p}} \lambda }{p}\]

OpenStudy (anonymous):

which must be equal to give F

OpenStudy (anonymous):

I don't know... your notation is rather confusing to me so I'm not quite sure what it says.